# A modern Fortran library for $\mathrm{SU}(3)$ coupling and recoupling coefficients

### v2

@@ Abstract
Keywords:
$\mathrm{SU}(3)$ coupling coefficients $\mathrm{SU}(3)$ recoupling coefficients
@@ 1 Introduction
Recently, in parallel with the development of the present library, a C++ implementation SU3lib of the DA algorithms has been developed by Dytrych et al. tdsu3lib . This library similarly provides for OpenMP multithreaded operation and supports the use of multiprecision arithmetic.
@@ 4.1 Module ndsu3lib_tools
Before a program first invokes ndsu3lib to calculate $\mathrm{SU}(3)$ coupling or recoupling coefficients, it must first initialize the library, by calling the subroutine initialize_ndsu3lib. To increase speed and avoid loss of precision, this subroutine allocates and recursively precalculates arrays containing binomial coefficients and, optionally, factors $I(i,j,k)$ and $S(i,j,k)$. These factors are defined in (104) and (109) and are needed only if $\mathrm{SU}(3)\supset\mathrm{SO}(3)$ RCCs are to be calculated. If the WIGXJPF library is being used, then initialize_ndsu3lib also initializes WIGXJPF. In OpenMP multithreaded applications, the subroutine initialize_ndsu3lib should be called separately by each thread.
@@ 5.1 $\mathrm{SU}(3)\supset\mathrm{U}(1)\times\mathrm{SU}(2)$ reduced coupling coefficients
Starting from $\Sigma_{w}=66$, the AD library produces incorrect results [^10] By incorrect results we mean an error greater than the greatest errors presented in this section by many ($\sim 10$) orders of magnitude., which is indicated by missing data in Fig. 2. Hence, ndsu3lib works for a larger range of quantum numbers. The precisions of the three libraries are comparable for small quantum numbers, however, with increasing quantum numbers the AD library and SU3lib lose precision more rapidly than ndsu3lib.
@@ 5.2 $\mathrm{SU}(3)\supset\mathrm{SO}(3)$ reduced coupling coefficients
Furthermore, WIGXJPF provides more reliable and accurate results than GSL for angular momentum coupling and recoupling coefficients, for large angular momenta wigxjpf . Usage of WIGXJPF can reduce the errors in the results calculated by ndsu3lib by several orders of magnitude compared to results obtained using GSL. Thus, we recommend using WIGXJPF as the library for angular momentum coupling and recoupling coefficients. All results shown in the present work are obtained using WIGXJPF.
Fig. 4 compares the maximal and mean errors of ndsu3lib, the AD library, and SU3lib as functions of $\Sigma_{w}$ (with quadruple precision arithmetic enabled for these other libraries, as well). For $\Sigma_{w}\geq 35$ (to the right of the gray vertical line), computations were made for only 100 randomly selected sets of $\mathrm{SU}(3)$ quantum numbers. With increasing quantum numbers, the errors tend to increase and then saturate (the mean error tends to slightly decrease). When the random sampling starts, the errors decrease little [^11] The decrease of the maximal error is not surprising because the random sampling is likely to exclude extremal cases. and start exhibiting less systematic behavior. For greater quantum numbers ndsu3lib is slightly more precise than the other two libraries. A systematic comparison for $\Sigma_{w}>46$ was not done due to very long computation time.
@@ 6.1 $\mathrm{SU}(3)\supset\mathrm{U}(1)\times\mathrm{SU}(2)$ reduced coupling coefficients
Figure 8: Time spent calculating the $\mathrm{SU}(3)\supset\mathrm{U}(1)\times\mathrm{SU}(2)$ RCCs divided by the number of $\mathrm{SU}(3)$ couplings as function of $\Sigma_{w}$ for ndsu3lib (squares), the AD library (circles), and SU3lib (crosses) (lower is better). Ratios of the time spent by ndsu3lib over the times spent by the AD library (circles) and SU3lib (crosses) are shown as well (lower is better for ndsu3lib).
@@ 6.2 $\mathrm{SU}(3)\supset\mathrm{SO}(3)$ coupling coefficients
Figure 9: Same as Fig. 8 but for $\mathrm{SU}(3)\supset\mathrm{SO}(3)$ RCCs.
@@ 6.3 $U$ recoupling coefficients
Figure 10: Time spent calculating the $U$ recoupling coefficients divided by the number of sets of the $\mathrm{SU}(3)$ quantum numbers in the sum $\Sigma_{r}$ as function of $\Sigma_{r}$ for ndsu3lib (squares), the AD library (circles), and SU3lib (crosses) (lower is better). Ratios of the time spent by ndsu3lib over the times spent by the AD library (circles) and SU3lib (crosses) are shown as well (lower is better for ndsu3lib).
@@ 6.4 $Z$ recoupling coefficients
Figure 11: Same as Fig. 10 but for $Z$ recoupling coefficients.
@@
Figure 12: Time spent calculating the 9-($\lambda,\mu$) coefficients divided by the number of sets of the $\mathrm{SU}(3)$ quantum numbers in the sum $\Sigma_{9}$ as function of $\Sigma_{9}$ for ndsu3lib (squares) and SU3lib (crosses) (lower is better). Ratio of the time spent by ndsu3lib and SU3lib is shown as well (lower is better for ndsu3lib).

### v3

@@ 1 Introduction
Recently, in parallel with the development of the present library, a C++ implementation SU3lib of the DA algorithms has been developed by Dytrych et al. tdsu3lib . This library similarly provides for OpenMP multithreaded operation and supports the use of multiprecision arithmetic. Since the original AD code, SU3lib, and ndsu3lib are all based on the same underlying DA algorithm, they yield consistent sets of coupling coefficients, namely, following the same phase conventions and the same prescriptions for the resolution of inner and outer multiplicities.
@@ 4 Structure, implementation details, and usage of the library
Since ndsu3lib implements the same underlying DA algorithm (section 3.1.3) as the AD and SU3lib libraries, all three libraries yield consistent sets of coupling (and recoupling) coefficients, differing only in matters of numerical precision (section 5). Namely, all three libraries follow the same phase convention (68) of Hecht hecht and the same BLH prescription (50) for the resolution of outer multiplicities in generating the canonical RCCs, from which the recoupling coefficients follow. Furthermore, in generating the $\mathrm{SU}(3)\supset\mathrm{SO}(3)$ reduced coupling coefficients, all three libraries follow the same prescription for resolution of the inner multiplicities (section 3.2), based on Gram-Schmidt orthonormalization (16) of the Elliott basis.
@@ 4.1 Module ndsu3lib_tools
Before a program first invokes ndsu3lib to calculate $\mathrm{SU}(3)$ coupling or recoupling coefficients, it must first initialize the library, by calling the subroutine initialize_ndsu3lib. To increase speed and avoid loss of precision, this subroutine allocates and recursively precalculates arrays containing binomial coefficients and, optionally, factors $I(i,j,k)$ and $S(i,j,k)$. These factors are defined in (104) and (109) and are needed only if $\mathrm{SU}(3)\supset\mathrm{SO}(3)$ RCCs are to be calculated. If the WIGXJPF library is being used, then initialize_ndsu3lib also initializes WIGXJPF. While ndsu3lib is safe to use in OpenMP multithreaded applications, the subroutine initialize_ndsu3lib should be called separately by each thread. In such applications the arrays of binomial coefficients and factors $I(i,j,k)$ and $S(i,j,k)$ are shared, and only one thread initializes them.
@@ 5.1 $\mathrm{SU}(3)\supset\mathrm{U}(1)\times\mathrm{SU}(2)$ reduced coupling coefficients
The precisions of the three libraries are comparable for small quantum numbers. However, with increasing quantum numbers the AD library and SU3lib lose precision more rapidly than ndsu3lib. Moreover, starting from $\Sigma_{w}=66$, there are examples of RCCs for which the AD library fails outright, by which we mean not simply an incremental loss of precision, but rather an error on the order of unity or greater. Consequently, no results are shown for the AD library in Fig. 2 for $\Sigma_{w}\geq 66$. In contrast, the orthonormality relation (34) was verified exhaustively for all sets of RCCs with $\Sigma_{w}\leq 81$ for ndsu3lib, with no such failures found, as can be seen from the maximal errors of $\lesssim 10^{-10}$ in Fig. 2 (top).
@@ 5.2 $\mathrm{SU}(3)\supset\mathrm{SO}(3)$ reduced coupling coefficients
Furthermore, WIGXJPF provides more reliable and accurate results than GSL for angular momentum coupling and recoupling coefficients, for large angular momenta wigxjpf . For large quantum numbers, use of WIGXJPF can reduce the errors in the results calculated by ndsu3lib by several orders of magnitude compared to results obtained using GSL. For example, for the coupling $(30,6)\times(5,10)\to(25,3)$, the maximal error obtained using GSL is approximately $10^{-11}$, whereas with WIGXJPF it is approximately $10^{-14}$. Thus, WIGXJPF should be used as the library for angular momentum coupling and recoupling coefficients, if maximal precision is required in calculations involving large values of quantum numbers. All results shown in the present work are obtained using WIGXJPF. However, use of GSL does not noticeably degrade the precision for the range of quantum numbers explored in the present study.
Fig. 4 compares the maximal and mean errors of ndsu3lib, the AD library, and SU3lib as functions of $\Sigma_{w}$ (with quadruple precision arithmetic enabled for these other libraries, as well). For $\Sigma_{w}\geq 35$ (to the right of the gray vertical line), computations were made for only 100 randomly selected sets of $\mathrm{SU}(3)$ quantum numbers. With increasing quantum numbers, the errors tend to increase and then saturate (the mean error tends to slightly decrease). When the random sampling starts, the errors decrease little [^10] The decrease of the maximal error is not surprising, because random sampling is likely to miss extremal cases. and start exhibiting less systematic behavior. For greater quantum numbers ndsu3lib is slightly more precise than the other two libraries. A systematic comparison for $\Sigma_{w}>46$ was not done due to very long computation time.
@@ 6.1 $\mathrm{SU}(3)\supset\mathrm{U}(1)\times\mathrm{SU}(2)$ reduced coupling coefficients
Figure 8: Time spent calculating $\mathrm{SU}(3)\supset\mathrm{U}(1)\times\mathrm{SU}(2)$ RCCs divided by the number of $\mathrm{SU}(3)$ couplings, as a function of $\Sigma_{w}$, for ndsu3lib (squares), the AD library (circles), and SU3lib (crosses) (lower is better). Ratios of the time spent by ndsu3lib over the times spent by the AD library (circles) and SU3lib (crosses) are shown as well (lower is better for ndsu3lib).
@@ 6.2 $\mathrm{SU}(3)\supset\mathrm{SO}(3)$ coupling coefficients
Figure 9: Time spent calculating $\mathrm{SU}(3)\supset\mathrm{SO}(3)$ RCCs, presented as described in the caption to Fig. 8.
@@ 6.3 $U$ recoupling coefficients
Figure 10: Time spent calculating $U$ recoupling coefficients divided by the number of sets of the $\mathrm{SU}(3)$ quantum numbers for the given value of $\Sigma_{r}$, as a function of $\Sigma_{r}$, for ndsu3lib (squares), the AD library (circles), and SU3lib (crosses) (lower is better). Ratios of the time spent by ndsu3lib over the times spent by the AD library (circles) and SU3lib (crosses) are shown as well (lower is better for ndsu3lib).
@@ 6.4 $Z$ recoupling coefficients
Figure 11: Time spent calculating $Z$ recoupling coefficients, presented as described in the caption to Fig. 10.
@@
Figure 12: Time spent calculating 9-($\lambda,\mu$) coefficients divided by the number of sets of the $\mathrm{SU}(3)$ quantum numbers for the given value of $\Sigma_{9}$, as a function of $\Sigma_{9}$, for ndsu3lib (squares) and SU3lib (crosses) (lower is better). Ratio of the time spent by ndsu3lib and SU3lib is shown as well (lower is better for ndsu3lib).